Diagraph of Metaphysic or Ontology

Jacob Lorhard, trans. by Sara L. Uckelman^{∗}

\begin{cases}Essence. See A. \\ Being. See B. \end{cases}

∗Many thanks to Brian M. Scott and Ursula Whitcher for help with one or two tricky bits, and Samson de Jager and Joel Uckelman for putting in nearly as much time making the TeX package doing this display as I did translating.